\begin{thebibliography}{59}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Atanasova et~al.(2023)Atanasova, Camburu, Lioma, Lukasiewicz,
  Simonsen, and Augenstein}]{atanasova2023faithfulness}
P~Atanasova, OM~Camburu, C~Lioma, T~Lukasiewicz, JG~Simonsen, and I~Augenstein.
  2023.
\newblock Faithfulness tests for natural language explanations.
\newblock In \emph{Proceedings of the Annual Meeting of the Association for
  Computational Linguistics}, volume~2, pages 283--294. Association for
  Computational Linguistics (ACL).

\bibitem[{Atanasova(2024)}]{atanasova2024diagnostic}
Pepa Atanasova. 2024.
\newblock A diagnostic study of explainability techniques for text
  classification.
\newblock In \emph{Accountable and Explainable Methods for Complex Reasoning
  over Text}, pages 155--187. Springer.

\bibitem[{Biran and Cotton(2017)}]{biran2017explanation}
Or~Biran and Courtenay Cotton. 2017.
\newblock Explanation and justification in machine learning: A survey.
\newblock In \emph{IJCAI-17 workshop on explainable AI (XAI)}, volume~8.

\bibitem[{Boyd-Graber et~al.(2022)Boyd-Graber, Carton, Feng, Liao, Lombrozo,
  Smith-Renner, and Tan}]{boyd-graber-etal-2022-human}
Jordan Boyd-Graber, Samuel Carton, Shi Feng, Q.~Vera Liao, Tania Lombrozo,
  Alison Smith-Renner, and Chenhao Tan. 2022.
\newblock \href {https://doi.org/10.18653/v1/2022.naacl-tutorials.4}
  {Human-centered evaluation of explanations}.
\newblock In \emph{Proceedings of the 2022 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies: Tutorial Abstracts}, pages 26--32, Seattle, United States.
  Association for Computational Linguistics.

\bibitem[{Camburu and Akata(2021)}]{camburu2021natural}
Oana-Maria Camburu and Zeynep Akata. 2021.
\newblock Natural-xai: Explainable ai with natural language explanations.
\newblock In \emph{International Conference on Machine Learning}.

\bibitem[{Camburu et~al.(2018)Camburu, Rockt{\"a}schel, Lukasiewicz, and
  Blunsom}]{camburu2018snli}
Oana-Maria Camburu, Tim Rockt{\"a}schel, Thomas Lukasiewicz, and Phil Blunsom.
  2018.
\newblock \href
  {http://papers.nips.cc/paper/8163-e-snli-natural-language-inference-with-natural-language-explanations.pdf}
  {e-snli: Natural language inference with natural language explanations}.
\newblock In \emph{Advances in Neural Information Processing Systems}, pages
  9539--9549. Curran Associates, Inc.

\bibitem[{Camburu et~al.(2020)Camburu, Shillingford, Minervini, Lukasiewicz,
  and Blunsom}]{camburu-etal-2020-make}
Oana-Maria Camburu, Brendan Shillingford, Pasquale Minervini, Thomas
  Lukasiewicz, and Phil Blunsom. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.382} {Make up your
  mind! adversarial generation of inconsistent natural language explanations}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 4157--4165, Online. Association for
  Computational Linguistics.

\bibitem[{Dalal et~al.(2024)Dalal, Valentino, Freitas, and
  Buitelaar}]{dalal-etal-2024-inference}
Dhairya Dalal, Marco Valentino, Andre Freitas, and Paul Buitelaar. 2024.
\newblock \href {https://doi.org/10.18653/v1/2024.acl-long.14} {Inference to
  the best explanation in large language models}.
\newblock In \emph{Proceedings of the 62nd Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 217--235,
  Bangkok, Thailand. Association for Computational Linguistics.

\bibitem[{Dalvi et~al.(2021)Dalvi, Jansen, Tafjord, Xie, Smith, Pipatanangkura,
  and Clark}]{dalvi2021explaining}
Bhavana Dalvi, Peter Jansen, Oyvind Tafjord, Zhengnan Xie, Hannah Smith,
  Leighanna Pipatanangkura, and Peter Clark. 2021.
\newblock Explaining answers with entailment trees.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pages 7358--7370.

\bibitem[{Deutsch(2011)}]{deutsch2011beginning}
David Deutsch. 2011.
\newblock \emph{The beginning of infinity: Explanations that transform the
  world}.
\newblock Penguin UK.

\bibitem[{Geirhos et~al.(2020)Geirhos, Jacobsen, Michaelis, Zemel, Brendel,
  Bethge, and Wichmann}]{geirhos2020shortcut}
Robert Geirhos, J{\"o}rn-Henrik Jacobsen, Claudio Michaelis, Richard Zemel,
  Wieland Brendel, Matthias Bethge, and Felix~A Wichmann. 2020.
\newblock Shortcut learning in deep neural networks.
\newblock \emph{Nature Machine Intelligence}, 2(11):665--673.

\bibitem[{Guidotti et~al.(2018)Guidotti, Monreale, Ruggieri, Turini, Giannotti,
  and Pedreschi}]{guidotti2018survey}
Riccardo Guidotti, Anna Monreale, Salvatore Ruggieri, Franco Turini, Fosca
  Giannotti, and Dino Pedreschi. 2018.
\newblock A survey of methods for explaining black box models.
\newblock \emph{ACM computing surveys (CSUR)}, 51(5):1--42.

\bibitem[{He et~al.(2024)He, Wu, Camburu, Minervini, and
  Stenetorp}]{he-etal-2024-using}
Xuanli He, Yuxiang Wu, Oana-Maria Camburu, Pasquale Minervini, and Pontus
  Stenetorp. 2024.
\newblock \href {https://doi.org/10.18653/v1/2024.acl-long.728} {Using natural
  language explanations to improve robustness of in-context learning}.
\newblock In \emph{Proceedings of the 62nd Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 13477--13499,
  Bangkok, Thailand. Association for Computational Linguistics.

\bibitem[{Hempel and Oppenheim(1948)}]{hempel1948studies}
Carl~G Hempel and Paul Oppenheim. 1948.
\newblock Studies in the logic of explanation.
\newblock \emph{Philosophy of science}, 15(2):135--175.

\bibitem[{Jansen et~al.(2016)Jansen, Balasubramanian, Surdeanu, and
  Clark}]{jansen2016s}
Peter Jansen, Niranjan Balasubramanian, Mihai Surdeanu, and Peter Clark. 2016.
\newblock What’s in an explanation? characterizing knowledge and inference
  requirements for elementary science exams.
\newblock In \emph{Proceedings of COLING 2016, the 26th International
  Conference on Computational Linguistics: Technical Papers}, pages 2956--2965.

\bibitem[{Jansen et~al.(2021)Jansen, Smith, Moreno, and
  Ortiz}]{jansen2021challenges}
Peter Jansen, Kelly~J Smith, Dan Moreno, and Huitzilin Ortiz. 2021.
\newblock On the challenges of evaluating compositional explanations in
  multi-hop inference: Relevance, completeness, and expert ratings.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pages 7529--7542.

\bibitem[{Jansen et~al.(2018)Jansen, Wainwright, Marmorstein, and
  Morrison}]{jansen2018worldtree}
Peter Jansen, Elizabeth Wainwright, Steven Marmorstein, and Clayton Morrison.
  2018.
\newblock \href {https://www.aclweb.org/anthology/L18-1433} {{W}orld{T}ree: A
  corpus of explanation graphs for elementary science questions supporting
  multi-hop inference}.
\newblock In \emph{Proceedings of the Eleventh International Conference on
  Language Resources and Evaluation ({LREC} 2018)}, Miyazaki, Japan. European
  Language Resources Association (ELRA).

\bibitem[{Khashabi et~al.(2019)Khashabi, Azer, Khot, Sabharwal, and
  Roth}]{khashabi2019capabilities}
Daniel Khashabi, Erfan~Sadeqi Azer, Tushar Khot, Ashish Sabharwal, and Dan
  Roth. 2019.
\newblock On the capabilities and limitations of reasoning for natural language
  understanding.
\newblock \emph{arXiv preprint arXiv:1901.02522}.

\bibitem[{Kitcher(1989)}]{kitcher1989explanatory}
Philip Kitcher. 1989.
\newblock Explanatory unification and the causal structure of the world.

\bibitem[{Lewis et~al.(2021)Lewis, Stenetorp, and Riedel}]{lewis2021question}
Patrick Lewis, Pontus Stenetorp, and Sebastian Riedel. 2021.
\newblock Question and answer test-train overlap in open-domain question
  answering datasets.
\newblock In \emph{Proceedings of the 16th Conference of the European Chapter
  of the Association for Computational Linguistics: Main Volume}, pages
  1000--1008.

\bibitem[{Lombrozo(2012)}]{lombrozo2012explanation}
Tania Lombrozo. 2012.
\newblock Explanation and abductive inference.
\newblock \emph{Oxford handbook of thinking and reasoning}, pages 260--276.

\bibitem[{Miller(2019)}]{miller2019explanation}
Tim Miller. 2019.
\newblock \href {https://doi.org/https://doi.org/10.1016/j.artint.2018.07.007}
  {Explanation in artificial intelligence: Insights from the social sciences}.
\newblock \emph{Artificial Intelligence}, 267:1--38.

\bibitem[{Mitchell et~al.(1986)Mitchell, Keller, and
  Kedar-Cabelli}]{mitchell1986explanation}
Tom~M Mitchell, Richard~M Keller, and Smadar~T Kedar-Cabelli. 1986.
\newblock Explanation-based generalization: A unifying view.
\newblock \emph{Machine learning}, 1(1):47--80.

\bibitem[{Parcalabescu and Frank(2024)}]{parcalabescu2024measuring}
Letitia Parcalabescu and Anette Frank. 2024.
\newblock On measuring faithfulness or self-consistency of natural language
  explanations.
\newblock In \emph{Proceedings of the 62nd Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 6048--6089.

\bibitem[{Quan et~al.(2024{\natexlab{a}})Quan, Valentino, Dennis, and
  Freitas}]{quan-etal-2024-enhancing}
Xin Quan, Marco Valentino, Louise Dennis, and Andre Freitas.
  2024{\natexlab{a}}.
\newblock \href {https://aclanthology.org/2024.eacl-long.1} {Enhancing ethical
  explanations of large language models through iterative symbolic refinement}.
\newblock In \emph{Proceedings of the 18th Conference of the European Chapter
  of the Association for Computational Linguistics (Volume 1: Long Papers)},
  pages 1--22, St. Julian{'}s, Malta. Association for Computational
  Linguistics.

\bibitem[{Quan et~al.(2024{\natexlab{b}})Quan, Valentino, Dennis, and
  Freitas}]{quan2024verification}
Xin Quan, Marco Valentino, Louise~A Dennis, and Andr{\'e} Freitas.
  2024{\natexlab{b}}.
\newblock Verification and refinement of natural language explanations through
  llm-symbolic theorem proving.
\newblock \emph{arXiv preprint arXiv:2405.01379}.

\bibitem[{Rajani et~al.(2019)Rajani, McCann, Xiong, and
  Socher}]{rajani2019explain}
Nazneen~Fatema Rajani, Bryan McCann, Caiming Xiong, and Richard Socher. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1487} {Explain yourself!
  leveraging language models for commonsense reasoning}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 4932--4942, Florence, Italy.
  Association for Computational Linguistics.

\bibitem[{Salmon(1984)}]{salmon1984scientific}
Wesley~C Salmon. 1984.
\newblock \emph{Scientific explanation and the causal structure of the world}.
\newblock Princeton University Press.

\bibitem[{Salmon(2006)}]{salmon2006four}
Wesley~C Salmon. 2006.
\newblock \emph{Four decades of scientific explanation}.
\newblock University of Pittsburgh press.

\bibitem[{Schlegel et~al.(2020)Schlegel, Valentino, Freitas, Nenadic, and
  Batista-Navarro}]{schlegel2020framework}
Viktor Schlegel, Marco Valentino, Andr{\'e} Freitas, Goran Nenadic, and
  Riza~Theresa Batista-Navarro. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.lrec-1.660} {A framework
  for evaluation of machine reading comprehension gold standards}.
\newblock In \emph{Proceedings of The 12th Language Resources and Evaluation
  Conference}, pages 5359--5369, Marseille, France. European Language Resources
  Association.

\bibitem[{Sinha et~al.(2021)Sinha, Parthasarathi, Pineau, and
  Williams}]{sinha2021unnatural}
Koustuv Sinha, Prasanna Parthasarathi, Joelle Pineau, and Adina Williams. 2021.
\newblock Unnatural language inference.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 7329--7346.

\bibitem[{Stacey et~al.(2022)Stacey, Belinkov, and Rei}]{stacey2021supervising}
Joe Stacey, Yonatan Belinkov, and Marek Rei. 2022.
\newblock Supervising model attention with human explanations for robust
  natural language inference.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial
  Intelligence}, volume~36, pages 11349--11357.

\bibitem[{Thagard and Litt(2008)}]{thagard2008models}
Paul Thagard and Abninder Litt. 2008.
\newblock Models of scientific explanation.
\newblock \emph{The Cambridge Handbook of Computational Psychology}, pages
  549--564.

\bibitem[{Thayaparan et~al.(2022)Thayaparan, Valentino, Ferreira, Rozanova, and
  Freitas}]{thayaparan2022diff}
Mokanarangan Thayaparan, Marco Valentino, Deborah Ferreira, Julia Rozanova, and
  Andr{\'e} Freitas. 2022.
\newblock Diff-explainer: Differentiable convex optimization for explainable
  multi-hop inference.
\newblock \emph{Transactions of the Association for Computational Linguistics},
  10:1103--1119.

\bibitem[{Thayaparan et~al.(2020)Thayaparan, Valentino, and
  Freitas}]{thayaparan2020survey}
Mokanarangan Thayaparan, Marco Valentino, and Andr{\'e} Freitas. 2020.
\newblock \href {http://arxiv.org/abs/2010.00389} {A survey on explainability
  in machine reading comprehension}.
\newblock \emph{arXiv preprint arXiv:2010.00389}.

\bibitem[{Thayaparan et~al.(2021{\natexlab{a}})Thayaparan, Valentino, and
  Freitas}]{thayaparan2021explainable}
Mokanarangan Thayaparan, Marco Valentino, and Andr{\'e} Freitas.
  2021{\natexlab{a}}.
\newblock Explainable inference over grounding-abstract chains for science
  questions.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  ACL-IJCNLP 2021}, pages 1--12.

\bibitem[{Thayaparan et~al.(2024)Thayaparan, Valentino, and
  Freitas}]{thayaparan2024differentiable}
Mokanarangan Thayaparan, Marco Valentino, and Andr{\'e} Freitas. 2024.
\newblock A differentiable integer linear programming solver for
  explanation-based natural language inference.
\newblock In \emph{Proceedings of the 2024 Joint International Conference on
  Computational Linguistics, Language Resources and Evaluation (LREC-COLING
  2024)}, pages 449--458.

\bibitem[{Thayaparan et~al.(2021{\natexlab{b}})Thayaparan, Valentino, Jansen,
  and Ustalov}]{thayaparan-etal-2021-textgraphs}
Mokanarangan Thayaparan, Marco Valentino, Peter Jansen, and Dmitry Ustalov.
  2021{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/2021.textgraphs-1.17}
  {{T}ext{G}raphs 2021 shared task on multi-hop inference for explanation
  regeneration}.
\newblock In \emph{Proceedings of the Fifteenth Workshop on Graph-Based Methods
  for Natural Language Processing (TextGraphs-15)}, pages 156--165, Mexico
  City, Mexico. Association for Computational Linguistics.

\bibitem[{Turpin et~al.(2024)Turpin, Michael, Perez, and
  Bowman}]{turpin2024language}
Miles Turpin, Julian Michael, Ethan Perez, and Samuel Bowman. 2024.
\newblock Language models don't always say what they think: unfaithful
  explanations in chain-of-thought prompting.
\newblock \emph{Advances in Neural Information Processing Systems}, 36.

\bibitem[{Valentino and Freitas(2024)}]{valentino2024nature}
Marco Valentino and Andr{\'e} Freitas. 2024.
\newblock On the nature of explanation: An epistemological-linguistic
  perspective for explanation-based natural language inference.
\newblock \emph{Philosophy \& Technology}, 37(3):88.

\bibitem[{Valentino et~al.(2021{\natexlab{a}})Valentino, Pratt-Hartmann, and
  Freitas}]{valentino2021natural}
Marco Valentino, Ian Pratt-Hartmann, and Andr{\'e} Freitas. 2021{\natexlab{a}}.
\newblock Do natural language explanations represent valid logical arguments?
  verifying entailment in explainable nli gold standards.
\newblock In \emph{Proceedings of the 14th International Conference on
  Computational Semantics (IWCS)}, pages 76--86.

\bibitem[{Valentino et~al.(2022{\natexlab{a}})Valentino, Thayaparan, Ferreira,
  and Freitas}]{valentino2022hybrid}
Marco Valentino, Mokanarangan Thayaparan, Deborah Ferreira, and Andr{\'e}
  Freitas. 2022{\natexlab{a}}.
\newblock Hybrid autoregressive inference for scalable multi-hop explanation
  regeneration.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial
  Intelligence}, volume~36, pages 11403--11411.

\bibitem[{Valentino et~al.(2021{\natexlab{b}})Valentino, Thayaparan, and
  Freitas}]{valentino2021unification}
Marco Valentino, Mokanarangan Thayaparan, and Andr{\'e} Freitas.
  2021{\natexlab{b}}.
\newblock \href {https://www.aclweb.org/anthology/2021.eacl-main.15}
  {Unification-based reconstruction of multi-hop explanations for science
  questions}.
\newblock In \emph{Proceedings of the 16th Conference of the European Chapter
  of the Association for Computational Linguistics: Main Volume}, pages
  200--211, Online. Association for Computational Linguistics.

\bibitem[{Valentino et~al.(2022{\natexlab{b}})Valentino, Thayaparan, and
  Freitas}]{valentino2022case}
Marco Valentino, Mokanarangan Thayaparan, and Andr{\'e} Freitas.
  2022{\natexlab{b}}.
\newblock Case-based abductive natural language inference.
\newblock In \emph{Proceedings of the 29th International Conference on
  Computational Linguistics}, pages 1556--1568.

\bibitem[{Wei et~al.(2022)Wei, Wang, Schuurmans, Bosma, Xia, Chi, Le, Zhou
  et~al.}]{wei2022chain}
Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed~Chi, Quoc~V
  Le, Denny Zhou, et~al. 2022.
\newblock Chain-of-thought prompting elicits reasoning in large language
  models.
\newblock \emph{Advances in neural information processing systems},
  35:24824--24837.

\bibitem[{Wiegreffe and Marasovic()}]{wiegreffe1teach}
Sarah Wiegreffe and Ana Marasovic.
\newblock Teach me to explain: A review of datasets for explainable natural
  language processing.
\newblock In \emph{Thirty-fifth Conference on Neural Information Processing
  Systems Datasets and Benchmarks Track (Round 1)}.

\bibitem[{Wiegreffe and Marasovi{\'c}(2021)}]{wiegreffe2021teach}
Sarah Wiegreffe and Ana Marasovi{\'c}. 2021.
\newblock \href {http://arxiv.org/abs/2102.12060} {Teach me to explain: A
  review of datasets for explainable nlp}.
\newblock \emph{arXiv preprint arXiv:2102.12060}.

\bibitem[{Woodward and Ross(2021)}]{woodwardscientific}
James Woodward and Lauren Ross. 2021.
\newblock {Scientific Explanation}.
\newblock In Edward~N. Zalta, editor, \emph{The {Stanford} Encyclopedia of
  Philosophy}, {S}ummer 2021 edition. Metaphysics Research Lab, Stanford
  University.

\bibitem[{Xie et~al.(2020)Xie, Thiem, Martin, Wainwright, Marmorstein, and
  Jansen}]{xie2020worldtree}
Zhengnan Xie, Sebastian Thiem, Jaycie Martin, Elizabeth Wainwright, Steven
  Marmorstein, and Peter Jansen. 2020.
\newblock \href {https://www.aclweb.org/anthology/2020.lrec-1.671} {Worldtree
  v2: A corpus of science-domain structured explanations and inference patterns
  supporting multi-hop inference}.
\newblock In \emph{Proceedings of the 12th Language Resources and Evaluation
  Conference}, pages 5456--5473, Marseille, France. European Language Resources
  Association.

\bibitem[{Yang et~al.(2018)Yang, Qi, Zhang, Bengio, Cohen, Salakhutdinov, and
  Manning}]{yang2018hotpotqa}
Zhilin Yang, Peng Qi, Saizheng Zhang, Yoshua Bengio, William Cohen, Ruslan
  Salakhutdinov, and Christopher~D Manning. 2018.
\newblock Hotpotqa: A dataset for diverse, explainable multi-hop question
  answering.
\newblock In \emph{Proceedings of the 2018 Conference on Empirical Methods in
  Natural Language Processing}, pages 2369--2380.

\bibitem[{Yao et~al.(2024)Yao, Yu, Zhao, Shafran, Griffiths, Cao, and
  Narasimhan}]{yao2024tree}
Shunyu Yao, Dian Yu, Jeffrey Zhao, Izhak Shafran, Tom Griffiths, Yuan Cao, and
  Karthik Narasimhan. 2024.
\newblock Tree of thoughts: Deliberate problem solving with large language
  models.
\newblock \emph{Advances in Neural Information Processing Systems}, 36.

\bibitem[{Zhang et~al.(2024{\natexlab{a}})Zhang, Carvalho, and
  Freitas}]{zhang-etal-2024-learning}
Yingji Zhang, Danilo Carvalho, and Andre Freitas. 2024{\natexlab{a}}.
\newblock \href {https://doi.org/10.18653/v1/2024.acl-long.116} {Learning
  disentangled semantic spaces of explanations via invertible neural networks}.
\newblock In \emph{Proceedings of the 62nd Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 2113--2134,
  Bangkok, Thailand. Association for Computational Linguistics.

\bibitem[{Zhang et~al.(2024{\natexlab{b}})Zhang, Carvalho, Valentino,
  Pratt-Hartmann, and Freitas}]{zhang-etal-2024-improving}
Yingji Zhang, Danilo Carvalho, Marco Valentino, Ian Pratt-Hartmann, and Andre
  Freitas. 2024{\natexlab{b}}.
\newblock \href {https://aclanthology.org/2024.findings-eacl.97} {Improving
  semantic control in discrete latent spaces with transformer quantized
  variational autoencoders}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  EACL 2024}, pages 1434--1450, St. Julian{'}s, Malta. Association for
  Computational Linguistics.

\bibitem[{Zhang et~al.(2023)Zhang, Carvalho, Pratt-Hartmann, and
  Freitas}]{zhang2023towards}
Yingji Zhang, Danilo~S Carvalho, Ian Pratt-Hartmann, and Andre Freitas. 2023.
\newblock Towards controllable natural language inference through lexical
  inference types.
\newblock \emph{arXiv preprint arXiv:2308.03581}.

\bibitem[{Zhang et~al.(2024{\natexlab{c}})Zhang, Valentino, Carvalho,
  Pratt-Hartmann, and Freitas}]{zhang-etal-2024-graph}
Yingji Zhang, Marco Valentino, Danilo Carvalho, Ian Pratt-Hartmann, and Andre
  Freitas. 2024{\natexlab{c}}.
\newblock \href {https://doi.org/10.18653/v1/2024.findings-naacl.32}
  {Graph-induced syntactic-semantic spaces in transformer-based variational
  {A}uto{E}ncoders}.
\newblock In \emph{Findings of the Association for Computational Linguistics:
  NAACL 2024}, pages 474--489, Mexico City, Mexico. Association for
  Computational Linguistics.

\bibitem[{Zhao et~al.(2024)Zhao, Chen, Yang, Liu, Deng, Cai, Wang, Yin, and
  Du}]{zhao2024explainability}
Haiyan Zhao, Hanjie Chen, Fan Yang, Ninghao Liu, Huiqi Deng, Hengyi Cai,
  Shuaiqiang Wang, Dawei Yin, and Mengnan Du. 2024.
\newblock Explainability for large language models: A survey.
\newblock \emph{ACM Transactions on Intelligent Systems and Technology},
  15(2):1--38.

\bibitem[{Zhao et~al.(2023)Zhao, Geva, Lin, Yasunaga, Madaan, and
  Yu}]{zhao-etal-2023-complex}
Wenting Zhao, Mor Geva, Bill~Yuchen Lin, Michihiro Yasunaga, Aman Madaan, and
  Tao Yu. 2023.
\newblock \href {https://doi.org/10.18653/v1/2023.acl-tutorials.2} {Complex
  reasoning in natural language}.
\newblock In \emph{Proceedings of the 61st Annual Meeting of the Association
  for Computational Linguistics (Volume 6: Tutorial Abstracts)}, pages 11--20,
  Toronto, Canada. Association for Computational Linguistics.

\bibitem[{Zheng et~al.(2023)Zheng, Mishra, Chen, Cheng, Chi, Le, and
  Zhou}]{zheng2023take}
Huaixiu~Steven Zheng, Swaroop Mishra, Xinyun Chen, Heng-Tze Cheng, Ed~H Chi,
  Quoc~V Le, and Denny Zhou. 2023.
\newblock Take a step back: Evoking reasoning via abstraction in large language
  models.
\newblock \emph{arXiv preprint arXiv:2310.06117}.

\bibitem[{Zhu et~al.(2024)Zhu, Chen, Ye, Lyu, Tan, Marasovic, and
  Wiegreffe}]{zhu-etal-2024-explanation}
Zining Zhu, Hanjie Chen, Xi~Ye, Qing Lyu, Chenhao Tan, Ana Marasovic, and Sarah
  Wiegreffe. 2024.
\newblock \href {https://doi.org/10.18653/v1/2024.naacl-tutorials.3}
  {Explanation in the era of large language models}.
\newblock In \emph{Proceedings of the 2024 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies (Volume 5: Tutorial Abstracts)}, pages 19--25, Mexico City,
  Mexico. Association for Computational Linguistics.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{acl}
\usepackage{color, soul}
\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{multirow}
\usepackage{latexsym}
\usepackage{booktabs}
\usepackage{courier}
\usepackage{amsmath}
\usepackage{subfig}
\usepackage{placeins}
\usepackage{tikz}
\usepackage{hyperref}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{bbm}
\usepackage[ruled,vlined]{algorithm2e}

\makeatletter
\newcommand{\printfnsymbol}[1]{%
  \textsuperscript{\@fnsymbol{#1}}%
}
\makeatother

\title{Reasoning with Natural Language Explanations}

\author{Marco Valentino$^{1}$, Andr\'e Freitas$^{1,2,3}$\\  $^{1}$Idiap Research Institute, Switzerland,\\
$^{2}$Department of Computer Science, University of Manchester, UK\\ 
$^{3}$National Biomarker Centre, CRUK-MI, University of Manchester, UK\\
{\tt first.last@idiap.ch} \\}

\begin{document}
\maketitle

\begin{abstract}
Explanation constitutes an archetypal feature of human rationality, underpinning learning and generalisation, and representing one of the media supporting scientific discovery and communication. Due to the importance of explanations in human reasoning, an increasing amount of research in Natural Language Inference (NLI) has started reconsidering the role that explanations play in learning and inference, attempting to build explanation-based NLI models that can effectively encode and use natural language explanations on downstream tasks. 
Research in explanation-based NLI, however, presents specific challenges and opportunities, as explanatory reasoning reflects aspects of both material and formal inference, making it a particularly rich setting to model and deliver complex reasoning.
In this tutorial, we provide a comprehensive introduction to the field of explanation-based NLI, grounding this discussion on the epistemological-linguistic foundations of explanations, systematically describing the main architectural trends and evaluation methodologies that can be used to build systems capable of explanatory reasoning\footnote{Tutorial website: \url{https://sites.google.com/view/reasoning-with-explanations}}.
\end{abstract}

 \section{Introduction}

Building systems that can understand and explain the world is a long-standing goal for \emph{Artificial Intelligence (AI)} \cite{miller2019explanation,mitchell1986explanation,thagard2008models}. The ability to explain, in fact, constitutes an archetypal feature of human rationality, underpinning communication, learning, and generalisation, as well as one of the mediums enabling scientific discovery and progress through the formulation of explanatory theories \cite{lombrozo2012explanation,salmon2006four,kitcher1989explanatory,deutsch2011beginning}.


Due to the importance of explanation in human reasoning, an increasing amount of work has started reconsidering the role that explanation plays in learning and inference with natural language \cite{camburu2018snli,yang2018hotpotqa,rajani2019explain,jansen2018worldtree}. In contrast to the existing end-to-end paradigm based on Deep Learning, explanation-based NLI focuses on developing and evaluating models that can address downstream tasks through the explicit construction of a \emph{natural language explanation} \cite{dalvi2021explaining,jansen2016s,wiegreffe2021teach,stacey2021supervising}. In this context, explanation is seen as a potential solution to mitigate some of the well-known limitations in neural-based NLI architectures \cite{thayaparan2020survey}, including the susceptibility to learning via shortcuts, the inability to generalise out-of-distribution, and the lack of interpretability \cite{guidotti2018survey,biran2017explanation,geirhos2020shortcut,lewis2021question,sinha2021unnatural,schlegel2020framework}.

Research in explanation-based NLI, however, presents several fundamental challenges \cite{valentino2024nature}. 
First, the applied methodologies are still poorly informed by theories and accounts of explanations \cite{salmon2006four,woodwardscientific}. This gap between theory and practice poses the risk of slowing down progress, missing the opportunity to formulate clearer hypotheses on the inferential properties of natural language explanations and define systematic evaluation methodologies \cite{camburu-etal-2020-make,jansen2021challenges,atanasova2024diagnostic}. 
Second, explanation-based NLI models still lack robustness, control, and scalability for real-world applications.
In particular, existing approaches suffer from several limitations when composing explanatory reasoning chains and performing abstraction for NLI in complex domains \cite{khashabi2019capabilities,valentino2022hybrid}.

In this tutorial, we will provide a comprehensive introduction to explanatory reasoning in the context of NLI, by systematically categorising and surveying explanation-supporting benchmarks, architectures, and research trends.
Specifically, we will present how the understanding of explanatory inference have evolved in recent years, together with the emerging methodological and modelling strategies. In parallel, we will attempt to provide an epistemological-linguistic characterisation of natural language explanations reviewing the main theoretical accounts \cite{valentino2024nature,salmon2006four} to derive a fresh perspective for future work in the field. 

\section{Description}
 
This section outlines the content of the tutorial.

\subsection{Epistemological-Linguistic Foundations}

One of the main objectives of the tutorial is to provide a theoretically grounded foundation for explanation-based NLI, investigating the notion of explanation as a language and inference scientific object of interest, from both an \emph{epistemological} and \emph{linguistic} perspectives \cite{valentino2024nature,salmon2006four,jansen2016s}. 

To this end, we will present a systematic survey of the contemporary discussion in Philosophy of Science around the notion of a scientific explanation, attempting to shed light on the nature and function of explanatory arguments and their constituting elements. Here, we will critically review the main accounts of explanations, including the deductive-nomological and inductive-statistical account \cite{hempel1948studies}, the notion of statistical relevance and the causal-mechanical model \cite{salmon1984scientific}, and the unificationist account \cite{kitcher1989explanatory}, aiming to elicit what it means to perform explanatory reasoning.
Following the survey, we will focus on grounding the theoretical accounts for explanation-based NLI, attempting to identify the main feature of explanatory arguments in existing corpora of natural language explanations \cite{jansen2016s,xie2020worldtree,jansen2018worldtree}.

\subsection{Resources \& Evaluation Methods for Explanation-Based NLI}

In order to build NLI models that can reason through the generation of natural language explanations it is necessary to develop systematic evaluation methodologies. To this end, The tutorial will review the main resources, benchmarks and metrics in the field \cite{wiegreffe1teach}. 

Depending on the nature of the NLI problem, an explanation can include pieces of evidence at different levels of abstraction \cite{thayaparan2020survey}. Traditionally, the field has been divided into \emph{extractive} and \emph{abstractive} tasks. In extractive NLI, the reasoning required for the explanations is derivable from the original problem formulation, where the correct decomposition of the problem contains all the necessary inference steps for the answer \cite{yang2018hotpotqa}. On the other hand, abstractive NLI tasks require going beyond the surface form of the problem, where an explanation needs to account for and cohere definitions, abstract relations, which are not immediately available from the original context  \cite{jansen2021challenges,thayaparan-etal-2021-textgraphs}.

In addition, the tutorial will review the main evaluation metrics adopted to assess the quality of natural language explanations. Evaluating the quality of explanations, in fact, is a challenging problem as it requires accounting for multiple concurrent properties. Different metrics have been proposed in the field, ranging from reference-based metrics designed to assess the alignment between automatically generated explanations and human-annotated explanations \cite{camburu2018snli,jansen2021challenges}, and reference-free metrics designed to evaluate additional dimensions such as faithfulness \cite{parcalabescu2024measuring,atanasova2023faithfulness}, robustness \cite{camburu-etal-2020-make}, logical validity \cite{quan2024verification,valentino2021natural}, and plausibility \cite{dalal-etal-2024-inference}. 

\subsection{Explanation-Based Learning \& Inference}

We review the key architectural patterns and modelling strategies for reasoning and learning over natural language explanations. In particular, we focus on the following paradigms:

\paragraph{Multi-Hop Reasoning \& Retrieval-Based Models.} The construction of explanations typically requires multi-hop reasoning -- i.e., the ability to compose multiple pieces of evidence to support the final answer \cite{dalvi2021explaining,xie2020worldtree}. Multi-hop reasoning has been largely studied in a retrieval settings, where, given an external knowledge base, the model is required to select, collect and link the relevant knowledge required to arrive at a final answer \cite{valentino2022hybrid,valentino2021unification,valentino2022case}. Here, we will review the main retrieval-based architectures for multi-hop reasoning and explanation, highlighting some of the inherent limitations of such paradigm, including the tension between semantic drift and efficiency \cite{khashabi2019capabilities}.

\paragraph{Natural Language Explanation Generation.} In parallel with retrieval approaches, NLI using generative models have been used for supporting explanatory inference \cite{camburu2018snli,rajani2019explain}. In this setting, early approaches leverage human-annotated natural language explanations for training generative models \cite{dalvi2021explaining}. Subsequently, the advent of Large Language Models (LLMs) has made it possible to elicit explanatory reasoning via specific prompting techniques and in-context learning \cite{wei2022chain,yao2024tree,zheng2023take,he-etal-2024-using}. Here, we review the main trends in the LLM-based generative paradigms, highlighting persisting limitations such as hallucinations and faithfulness \cite{turpin2024language}.

\subsection{Semantic Control for Explanatory Reasoning}

Controlling the explanation generation process in neural-based models is particularly critical while modelling complex reasoning tasks. In this tutorial, we will review emerging trends which combine neural and symbolic approaches to improve semantic control in the explanatory reasoning process, which can provide formal guarantees on the quality of the explanations. These methods aim to integrate the content flexibility of language models (instrumental for supporting material inferences) and a formal inference properties.

In particular, we focus on the following key methods:

\paragraph{Leveraging Explanatory Inference Patterns for Explanation-Based NLI.} Inference patterns in explanation corpora can be leveraged to improve the efficiency and robustness of neural representations \cite{valentino2024nature,zhang2023towards}. In particular, we will review approaches that attempt to leverage the notion of unification power in corpora of natural language explanations to improve multi-hop reasoning in a retrieval setting and alleviate semantic drift \cite{valentino2022hybrid,valentino2021unification,valentino2022case}.

\paragraph{Constraint-Based Optimisation for Explanation-Based NLI.} We will focus on describing neuro-symbolic methods which target encoding explicit assumptions about the structure of natural language explanations \cite{thayaparan2021explainable}. Here, we will review methods performing multi-hop inference via constrained optimisation, integrating neural representations with explicit constraints via end-to-end differentiable optimisation approaches \cite{thayaparan2022diff,thayaparan2024differentiable}.

\paragraph{Formal-Geometric Inference Controls over Latent Spaces.} Covers emerging methodologies which focus on learning latent spaces with better representational properties for explanatory NLI, using language Variational Autoencoders (VAEs) for delivering better disentanglement and separability of language and inference properties \cite{zhang-etal-2024-learning,zhang-etal-2024-graph,zhang-etal-2024-improving,zhang-etal-2024-learning} which support better inference control. These methods deliver an additional geometrical structure to latent spaces, aiming to deliver the vision of 'inference as latent geometry'.

\paragraph{LLM-Symbolic Architectures} Finally, we will focus on hybrid neuro-symbolic architectures that attempt to leverage the material/content-based inference properties of LLMs for explanation generation with external symbolic approaches, which accounts for formal/logical validity refinement properties. In particular, we will review approaches that perform explanation refinement via the integration of LLMs and Theorem Provers to verify logical validity \cite{quan2024verification,quan-etal-2024-enhancing} and additional external tools to evaluate explanation properties such as uncertainty, plausibility and coherence \cite{dalal-etal-2024-inference}.

\section{Schedule}

The tutorial will be organised according to the following timeline:

\begin{enumerate}
    \item Introduction \& Motivation (20 min.)
    \item Epistemological-Linguistic Foundations (20 min.)
    \item Resources \& Evaluation for Explanation-Based NLI (40 min.)
    \item Explanation-Based Learning \& Inference (40 min.)
    \item Semantic Control for Explanatory Reasoning (40 min.)
    \item Synthesis, Discussion, and Q\&A (20 min.)
\end{enumerate}

 \section{Breadth \& Diversity}
 

The tutorial will cover a wide spectrum of topics in different fields, ranging from Philosophy, Machine Learning, Natural Language Processing, Knowledge Representation and Automated Reasoning. This diversity of topics will help create a rich environment in which academics from different backgrounds and cultural contexts can integrate different perspectives. The tutorial plan includes integrated open Q\&A sessions and practical demonstrations.
  
 \section{Prerequisites}
 
We do not expect attendees to be familiar with previous research on NLI and Explanatory inference. On the opposite, we intent this tutorial to be an efficient and deep onboarding into the state-of-the-art in those areas. Participants should have a general background knowledge in deep learning, including recent trends and architectures such as Large Language Models. Participants are expected to be familiar with some of the broader NLI tasks, such as Textual Entailment and Question Answering.
 
 \section{Reading List}

 \paragraph{Epistemological-Linguistic Foundations}

\paragraph{\citet{valentino2024nature}} On the Nature of Explanation: An Epistemological-Linguistic Perspective for Explanation-Based Natural Language Inference.

\paragraph{\citet{salmon2006four}} Four Decades of Scientific Explanation.

\paragraph{\citet{jansen2016s}} What’s in an Explanation? Characterizing Knowledge and Inference Requirements for Elementary Science Exam.

\paragraph{Resources, Models and Evaluation}
\paragraph{\citet{wiegreffe2021teach}} Teach me to Explain: A Review of Datasets for Explainable Natural Language Processing.
\paragraph{\citet{thayaparan2020survey}} A Survey on Explainability in Machine Reading Comprehension.
\paragraph{\citet{zhao2024explainability}} Explainability for Large Language Models: A Survey.

\paragraph{Related Tutorials}
\paragraph{\citet{zhu-etal-2024-explanation}} Explanation in the Era of Large Language Models.
\paragraph{\citet{camburu2021natural}} Natural-XAI: Explainable AI with Natural Language Explanation.
\paragraph{\citet{zhao-etal-2023-complex}} Complex Reasoning in Natural Language.
\paragraph{\citet{boyd-graber-etal-2022-human}} Human-Centered Evaluation of Explanations.

 \section{Instructor information}
 \paragraph{Marco Valentino, {\normalfont{Idiap Research Institute}}.\protect\footnote{\url{mailto:marco.valentino@idiap.ch}}} Marco is a postdoctoral researcher at the Idiap Research Institute, Switzerland. His research is carried out at the intersection of Natural Language Inference and Neuro-Symbolic models focusing on building systems that can reason through natural language explanations in complex domains (e.g., mathematics, science, biomedical and clinical applications, ethical reasoning). He has published papers in major AI and NLP conferences including AAAI, ACL, EMNLP, NAACL and EACL. Marco was involved in the organisation of workshops including MathNLP (EMNLP 2022 and LREC-COLING 2024), and TextGraphs (COLING 2022 and ACL 2024).

 \paragraph{Andr\'e Freitas, {\normalfont{University of Manchester \& Idiap Research Institute}}.\protect\footnote{\url{mailto:andre.freitas@manchester.ac.uk}}} André Freitas leads the Neuro-symbolic AI Lab at the University of Manchester and IDIAP Research Institute. His main research interests are on enabling the development of AI methods to support abstract, flexible and controlled reasoning in order to support AI-augmented scientific discovery. In particular, he investigates how the combination of neural and symbolic data representation paradigms can deliver better models of inference. He is an active contributor to the main conferences and journals in the AI/Natural Language Processing (NLP) interface (AAAI, NeurIPs, ACL, EMNLP, EACL, COLING, TACL, Computational Linguistics), with over 100 peer-reviewed publications. He contributed to the organisation of MathNLP at EMNLP 2022 and LREC-COLING 2024. Andr\'e participated in 7 tutorials, and co-organised 1 conference and 6 workshops.

\section*{Acknowledgements}
This work was partially funded by the Swiss National Science Foundation (SNSF) project NeuMath (\href{https://data.snf.ch/grants/grant/204617}{200021\_204617}).

\bibliography{acl}
\bibliographystyle{acl_natbib}

\end{document}